OpenStudy (anonymous):

the function h is defined by h(x)=3x^2-2. Find h(5z)

OpenStudy (anonymous):

simply plug in x=5z so it will be 3(5z)^2-2 =3*25z^2-2=75z^2-2

OpenStudy (anonymous):

Thank you. And what about h(x)=3x^2+2x h(4x)

OpenStudy (anonymous):

Would I just do 3(4x)^2+2(4x)?

OpenStudy (anonymous):

yes exactly...

OpenStudy (anonymous):

So i would get 12x^2+8x

OpenStudy (anonymous):

no...4^2=16 so 16*3=48x^2+8x

OpenStudy (anonymous):

lousy mistake...THANK YOU SO MUCH